A proton moves from point A to point B. Its distance from point A is 2mm and from point B is 4 mm. Find: a. the potential difference from point A to B b. work done on the proton


1
Expert's answer
2021-05-24T16:22:48-0400

Gives

Point P is another point

AP=2mm

BP=4mm

q="1.6\times10^{-19}c"

Particle is Proton

Due to poin A potential difference

"V_A=\frac{kq}{r^2}"

"V_A=\frac{9\times10^{9}\times1.6\times10^{-19}}{2\times10^{-3}}"

"V_A=7.2\times10^{-7}V"

"V_B=\frac{kq}{r^2}"

"V_B=\frac{9\times10^9\times1.6\times10^{-19}}{4\times10^{-3}}"

"V_B=3.6\times10^{-7}V"

Potential difference between A to B"V_{AB}=V_A-V_B"

"V_{AB}={7.2 \times10^{-7}}-{3.6\times10^{-7}}"

"V_{AB}=3.6\times10^{-7}V"

Part (b)

Work done by proton

"W=q\u2206V"

"W=1.6\times10^{-19}\times3.6\times10^{-7}"

"W=5.76\times10^{-26}J"